In 7 years, rica's age will exceed BC's age by 14 years. Last year, the sum of their ages was 20. how old is Rica?
Answer by THANApHD(104)

let rica's present age is x
let BC's present age is y

so after 7 years
x+7 = y+7+14
x-y = 14 -----------1
last year
x-1+y-1=20
x+y = 22 ----------2

1+2 => 2x = 36
x=18
y=4